# Kuan-Cheng Chen QuEST Imperial College London London, United Kingdom

### v2

@@ Acknowledgment
    The authors extend their gratitude to W.C. Qian for invaluable assistance and insights, which have been pivotal to the success of this research work. K.C. is grateful for the financial support from both the Turing Scheme for the Imperial Global Fellows Fund and the Taiwanese Government Scholarship to Study Abroad. This work was supported by the National Science and Technology Council (NSTC), Taiwan, under Grant NSTC 112-2119-M-007-008- and NSTC 113-2119-M-007-013-. The authors would like to thank the National Center for High-performance Computing of Taiwan for providing computational and storage resources. Additionally, the authors would like to extend special thanks to the technical support provided by the NVAITC and NVIDIA Quantum team.
    Code Visability
    The source code for generating the dataset and figures presented in this research work is openly accessible at https://github.com/Tim-Li/cuTN-QSVM.git.

### v3

@@ Acknowledgment
    The authors express their gratitude to W.C. Qian (MediaTek) for invaluable assistance and insights, which were pivotal to the success of this research. This research was funded by the U.K. Engineering and Physical Sciences Research Council under Grant No. EP/W032643/1. K.C. acknowledges financial support from the Turing Scheme for the Imperial Global Fellows Fund and the Taiwanese Government Scholarship to Study Abroad. This work was also supported by the National Science and Technology Council (NSTC), Taiwan, under Grants NSTC 112-2119-M-007-008- and NSTC 113-2119-M-007-013-. The authors thank the National Center for High-performance Computing of Taiwan for providing computational and storage resources, as well as the NVAITC and NVIDIA Quantum team for their technical support.